OpenStudy (anonymous):

please help me with this problem , i need to know what the steps are on how to find x ---- 28-6(2x+4)=124 .

hartnn (hartnn):

first use the distributive property \(a(b+c)=ab+ac\)

hartnn (hartnn):

and tell me what u get for 6(2x+4)

OpenStudy (cathyangs):

Is that really the best course of action? I would think that subtracting 28, and then dividing by -6...etc would be faster.

OpenStudy (anonymous):

yeah i dont remember the teacher saying anything about doing that hartnn

hartnn (hartnn):

this action will lead to dividing by 12 in the end,whicj is simpler than dividing by -6

OpenStudy (anonymous):

im confused can somone write out all the steps , then i can study what you did ? i have other problems like this on my homework i need to know how to solve this type of problem

hartnn (hartnn):

i mean dividing by -6 and 2 also *

OpenStudy (cathyangs):

Personally, this is my approach: 28-6(2x+4)=124 ---Subtract 28 -6(2x+4)=96 ---multiply by -1, then divide by 6 (2x+4)=-96/6=-16 ---rewrite as 2x+4=-16 ---subtract 4 2x=-20 ---divide by 2 x=-10 This is my way of doing these kinds of questions, but there are definitely others.

OpenStudy (anonymous):

i like this way , it seems easy to understand thank you! !

OpenStudy (cathyangs):

You're welcome! :) I'm glad to help.

OpenStudy (cwrw238):

28-6(2x+4)=124 distribute 28 - 12x - 24 = 124 28 - 24 - 124 = 12x -120 = 12x x = -10

OpenStudy (cwrw238):

either way is same amount of work

OpenStudy (anonymous):

acually that way of doing it looks more familiar